If there price of apples is R5 each and the price of Orange is R10 each? Explain what the absolute and relative prices of these fruits are


1
Expert's answer
2021-05-13T10:29:01-0400

Case of Apple 

Absolute Price = R5 

"Relative \space Price = \frac{Price \space of \space Given \space Product\space (Apple)}{ Price \space of\space Other\space Good\space (Orange)}"


"Thus \space Pr = \frac{R5}{R10}\space i.e \space 1:2"

 

Case of orange

Absolute Price = R10

"Relative\space Price =\frac{Price \space of\space Given\space Product\space (Orange)}{ Price\space of\space Other\space Good\space (Apple)}"


"Pr= \frac{R10}{R5}\space i.e \space 2:1"